Conversion of Star Wars to Adventure!

Funksaw

December 2002

Abstract

This is my partial conversion of Star Wars to Adventure. I also have a fuller conversion of Star Wars to a system extremely similar to Adventure, but since I don’t own the rights to Star Wars, nor is the house system dissimilar enough to Adventure to completely distinguish it in the eyes of the law. In other words: Not quite yet, and only to my close friends.

1 The Force and Inspiration—one and the same.

Mechanically, the Force is extremely similar to Inspiration, yet subtly different. For Star Wars, there are two force bars of Permanent and Temporary marks—“Light Side” and “Dark Side.” Light side governs the Reflective and Intuitive facets—Dark Side covers the Destructive facet. As a general rule, Daredevils tend to have higher Intuitive facets than Reflective facets, and Jedi tend to have higher Reflective facets than Intuitive facets. Characters cannot start with a Destructive Facet rating, as they should not start off with any permanent Dark Side points.

Characters receive 1 dot in Reflective or Intuitive facets for every light side force point they have.

“The Force is what gives a Jedi his power. It’s an energy field created by all living things. It binds the galaxy together.”

Obi-Wan Kenobi

A main character (PC or significant NPC) can use the Force, as it flows through all living things. Only Jedi, however, can use the Force in a conscious manner. However, even normal people can be pushed to their limits, and even beyond it. That is when the unconsciously use the Force.

The Force can be used for a number of things, as a game mechanic. First off, it is the fuel for most Jedi powers. However, it can be used to increase someone’s chances at succeeding at an action, or even to save the characters neck via a “deus ex Force” when it really matters. In short, you may spend force points on the following:

• Activating a Jedi Power (“Impressive. Most Impressive.”)

• Dramatic Editing (“I keep these panels for smuggling. I never thought I’d be smuggling myself in them!”)

• Cliffhanger (“Threepio! Tell Artoo to shut down all the garbage smashers on the detention level!”)

• Double a dicepool (“Great shot kid! That was one in a million!”)

• Get a hint about how to proceed. (“Chewbacca here is first mate on a vessel that may suit our purposes.”)

Force points may be regained in the following ways:

• Meditation. A character that relaxes for at least 2 days will replenish a force point, a Jedi that meditates can accomplish the same in 8 hours.

• Life-Affirming Heroism. Since the Force is the energy of all living things, a heroic action which results in the saving of life typically results in some sort of Force replenishment for all involved.

• Wild Success. If a character receives more than 5 successes on a roll (but gets no benefit from such an amazing result) the character receives a force point from self confidence.

The Dark Side

Yoda: “A Jedi’s strength flows from the Force. But beware of the Dark Side. Anger. Fear. Agression. The Dark Side of the Force are they. Easily they flow, quick to join you in a fight. If once you start down the Dark path, forever will it dominate your destiny, consume you it will, as it did Obi-Wan’s apprentice.” Luke: “Vader. Is the Dark Side stronger?” Yoda: “No… no… no. Quicker, easier, more seductive.” Luke: “But how amd I to know the good side from the bad?” Yoda: “You will know. When you are calm, at peace. Passive. A Jedi uses the Force for knowledge and defense, never for attack.”

That is the rub, of course. Unlike “normal” force points, Dark Side points accumulate, rather than subtract. Whenever the character spends Force points, as he normally would to accomplish a task counter to the Force, (such as murder in cold blood, eating a kitten, or something along those lines,) or uses the Force out of Anger, the character does not lose force points. Rather, they gain temporary dark force points equal to the amount of Force that they would have spent. When this pool surpasses the current Dark Side permanent rating, a Permanent Dark Side Point is added, and the temporary points are reset. (In other words, the first dark side point is gained when the character uses 1 dark point, the second at 2, the third at 3, etc.)

It is possible to use your Dark Force as a “bottomless pool” of Force points. Any character can get angry enough to resolve his way out of a situation through blind rage. This, however, is a good way to rack up a goodly sized pool of Dark points.

For each Dark Side Point, the character gains a dot in the Destructive Facet—this is a benefit of using the Dark Side to achieve your ends—these points do not go away if the Dark Side points are bought off via experience points. Secondly, the character gains a “dark vice” determined by the GM: Greed, paranoia, avarice, Impatience, etc. that in some way relates to the way the Force was used. Whenever the character comes across a situation that relates to the Dark Force flaw, the character must spend a Willpower point to avoid taking that action.

Dark Side points may be bought off with experience points—and with roleplaying—at a rate equal to what it would cost to gain that dot if it were a Light Side point. The deeper you dig yourself into the pit, the harder it is to get out.

A character with more permanent Dark Side points than permanent Light Side points has fallen to the Dark Side. It is up to individual GMs to determine if the character can still be played, or if it becomes an NPC.

Redemption: Characters with less than 10 Dark Side points can potentially be redeemed through roleplay and experience. It is a tough road though, and most do not have the strength of will to redeem fully.

1.1 Knacks, Jedi Powers, and Racial Skills

As a general rule, Knacks (Heroic Knacks) belong to Daredevils, Jedi Powers (Psychic Knacks) belong to Jedi, and Racial Skills (Dynamic Knacks) belong to Aliens. But there are some exceptions to the rules presented in Adventure which will be listed here:

1. Any character may take Daredevil Knacks, including Jedi and Aliens. Obi-Wan’s a good pilot, for example. This means Daredevils need to struggle to protect their schtick.

2. Fists of Stone and Lighting Reflexes may be bought as a Racial Skills

3. Telluric Resonance is known as “Force Sensitivity” for the purpose s of the setting.

4. Flame/Frost conjuration can also be used to throw lighting from your hands.

5. Psychic Synergy may also be used to communicate “tele-empathically” with another character. Exact words are not possible, but the person will automatically know where you are and what you need as an emotional feeling. The target must be willing and at least nominally allied with you, however. (This is the Jedi power Luke uses when contacting Leia while she’s driving the Millenium Falcon with Lando and Chewie as he’ s dangling from Cloud City.)

6. A Single Bound can be purchased as a Jedi Power.

2 New Backgrounds:

2.1 Droids

Droids are the number of mechanical servants you have. Droids never have dicepools, and are not creatures of the force—they are NPC s that are mechanically perfect in one or a number of areas, but usually incompetent in most others. For example, a protocol droid may know nothing about combat, while a combat droid may not even have the ability to speak. Droids can also help out on certain types of rolls—an Astromech droid, for example, may be able to boost the reliability of a fighter or other ship. The major types of droids are:

• Repair Astromech (R)

• Conversation & Protocol (C)

• Battle Droids (B)

• Medical Droids (Me)

• Imperial Droids (I)

• Maintenance Droids (Ma)

• Power Droids (P)

Several droids can perform the function of other droids, but are usually not as adept at it—the general rule is that the background rating equals the number of functions your droids, in total, have. A rating of 4 could represent 4 specialized droids, a single droid with four functions, or two dual-function droids, for example.

2.2 Fleet

Fleet determines how good your ship is: For every dot in Fleet, you receive 5 points to place into the ship’s statistics (See Space Combat) To upgrade the class of your ship, 10 points must be spent. The fleet background can exceed 6, and can be bought many times.

2.2.1 Space Combat

Ships come in three main classes: Fighter (Small), Cruiser (Medium), and Flagship (Large.) Mechanically, this has a number of effects.

When fighting against ships of different classes, the successes achieved on rolls are modified. Larger ships have more battery power—thus more powerful lazers, smaller ships are deft and maneuverable.

When a smaller ship fights a bigger ship, the following happens: The smaller ship doubles it’s successes on maneuver rolls related to dodging fire—smaller targets are

harder to hit. The smaller ship halves it’s successes on firepower damage, and the larger ship doubles it’s firepower damage.

A ship fighting a class two sizes larger than itself finds these effects doubled. Ships have 7 main statistics, outlined below, and a number of secondary statistics. Each point placed

into these statistics costs one of the “fleet” background points as above, and are usually paired with a skill.

Speed: Speed is how fast the ship can travel in sublight speeds. Useful for outrunning slow and bulky pursuers. Usually used in conjuction with “Pilot.” A ship without speed travels very slowly indeed, and cannot hope to outrun pursuers or dodge attacks.

Shielding: Shielding determines a number of things. First off, it acts as a “soak” for your craft. Secondly, a number of “Shields” healt h levels are added to your craft equal to your Shielding rating. As these health levels are destroyed, the shield score also goes down for “soaking.” However, shield levels replenish at a rate of 1 level per round.

Maneuver: Maneuver is how quick and agile the ship is—as opposed to it’s raw speed. Maneuver is usually used in conjunction with Pilot to dodge asteroids and lazerbolts.

Firepower: Firepower determines the number of dice added to a successful damage roll’s damage result. A ship that is unmanueverable must come to a full stop before turning.

Hyperdrive: Hyperdrive is a measure of how quickly the ship can travel in hyperspace. Useful mostly for long-range pursuits, as well as determining how much lead time you have in evading a persistent pursuer.

Navacomputer: Navacomputer is a measure of how quickly the Navacomputer can calculate the nessessary equations to posit a path through hyperspace that doesn’t involve you materializing near a black hole or in the middle of a supernova. And that’d end your trip real quick, kid. The Navacomputer score is rolled, once per turn. 5 successes are needed (cumulative) before the Navacomputer is finished with it’s calculations.

Reliability: When an attack gets through the ships shields, and the ship begins to take actual damage, the reliability score is rolled. If no successes are rolled on the reliability roll (or the ship has no reliability) a malfunction occurs. Secondly, reliability adds a number of health levels onto your craft equal to the reliability score. However, each one of those health levels has a penalty, -1 to all actions for the first one, -2 for the second one, etc. Once a reliability level has been damaged, the damage remains until the ship can be repaired.

2.2.2 Malfunctions

When a ship fails a reliability check, roll 2d10 and take the lowest of the two numbers and consult this chart:

1. Your hyperdrive is knocked out.

2. If you have an astromech droid, the shot damaged the droid to the point where it will need significant repair at a dock.

3. The blast caught an engine on the left or right, and you’re spinning You may only turn in one direction. Basic: Maneuverability is −2

4. The blast caught a weapons battery, and your weapons aren’t as effective, your damage rating is halved.

5. The blast has sent you into a tailspin—you continue on your course for three rounds, unable to turn. You cannot use Maneuverability to dodge

6. Your shields go completely down until they can be fixed by hand or by an astromech droid. It usually takes about 3 rounds—less a round for each droid or person assisting in the repair, and at least one round.

7. The blast has set parts of your ship on fire. Don’t ask me how thing s burn in space—what am I? Einstein? Anyway, if you can’t find a way to put it out in three rounds, you will take another level of damage as the fuel tanks explode—which also places your speed rating at a maximum o f 1.

8. The blast sent something flying in the cockpit—anyone within the ship takes a level of lethal damage.

9. Life support begins to fail. You’ve got about 4 minutes of oxygen before you being to get dizzy, and about 10 minutes of oxygen before you suffocate.

10. The blast went through your ship like sh*t through a bantha— barring the miracles of the force, you’ve got about 30 seconds as the ship explodes and whoever isn’t incinerated is exposed to hard vacuum.

3 Weapons and Armor and Other Goodies

Blaster: Acc 0, Dam 3L, Str 1 Min, Range 50, Rof: 3, Conc: P, Cost * Blaster Rifle: Acc +2, Dam 7L, Str 2 Min, Range 200, Rof:1 Conc: P, Cost ** Thermal Detonator: Damage 9L, Full damage to all targets within 10 meters, half damage to targets

within 20 meters.

Stun Weapon: Acc -1, Dam 9B, Str 1 Min, Range 20, Rof 1, Conc J, Cost * * Lightsaber: Gadget Rating 1, Dam Str+5A, Melee, Conceal J, Cost *** Blast Padding: 2L, 4B Cost * Flight Helmet: 5L, (called shots to head) Cost* Stormtrooper Armor: 3L, 5B Cost ** Exoskeleton, like Darth Vader’s: 6L, 9B Cost ****

A Quotes to show you how applicable certain knacks are to Star Wars

Dramatic Entrance: “Search the vessel and bring me those plans” —Vader Eagle Eyes: “That’s not a moon. That’s a space station.” —Ben Kenobi. Fists of Stone: “A droid don’t tear your arms off when they lose. Wookies have been known to do that.”

—Han Solo Gadgeteer: “But I was going over to Toshi Station to pick up some power converters!” —Luke Indomintable Will: “You weak minded fool! He’s using an old Jedi Mind Trick!” Instant Expert: “Your father wanted me to give this to you when you came of age… newwomm… bwwwoomm… ”

—Ben Kenobi Jack of all Tounges: “Uta Greeba Solo?” —Greedo. (WHO NEVER FIRED!) Lightning Reflexes: “That was too close” —Lando. Master of Dissimulation: “Negative, we’ve got a… um… reactor leak, very dangerous, just give us a few

minutes to log it down.” —Han Solo.

One Man Army: “Bring ’em on, I prefer a straight fight to all this sneaking around!” —Han Solo. Trick Shot: “Shoot the door, Kid!” —Han Solo Marked Man: “Skywalker is on that planet. I can sense it.” —Vader Force Sensitivity: “I have felt the presence of my old master. He is here.” —Vader Inspirational Aura: “I’m not sure, sir. they seem to believe that I am thier God.” —Threepio Psychic Control: “You don’t need to see his identification. These aren’t the droids you’re looking for.”

—Ben Kenobi.